OpenStudy (anonymous):

if f(x+y)=f(x).f(y) forr all real x,y and f'(0)=3,f(5)=2 then how can find f'(5)???

OpenStudy (michele_laino):

we have: \[\begin{gathered} f(0) = f\left( {0 + 0} \right) = {\left( {f(0)} \right)^2} \hfill \\ f\left( 0 \right)\left( {1 - f\left( 0 \right)} \right) = 0 \hfill \\ f(0) = 1 \hfill \\ f(0) = 0 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

then??

OpenStudy (anonymous):

Setting \(y=0\) gives \[f(x+0)=f(x)f(0)~~\iff~~f(x)=f(x)f(0)\] and differentiating gives \[f'(x)=f'(x)f(0)\] for all \(x\). This would suggest that \(f(0)=1\), since we know that \(f'(0)\neq0\).

OpenStudy (michele_laino):

we have: \[\begin{gathered} \frac{{f\left( {h + 5} \right) - f\left( 5 \right)}}{h} = \frac{{f\left( h \right)f\left( 5 \right) - f\left( 5 \right)}}{h} = f\left( 5 \right)\frac{{f\left( h \right) - 1}}{h} \hfill \\ \frac{{f\left( h \right) - f\left( 0 \right)}}{h} = \frac{{f\left( {h + 0} \right) - f\left( 0 \right)}}{h} = f\left( 0 \right)\frac{{f\left( h \right) - 1}}{h} \hfill \\ f'\left( 5 \right) = \mathop {\lim }\limits_{h \to 0} \frac{{f\left( {h + 5} \right) - f\left( 5 \right)}}{h} = \mathop {\lim }\limits_{h \to 0} \frac{{f\left( h \right)f\left( 5 \right) - f\left( 5 \right)}}{h} = f\left( 5 \right)\mathop {\lim }\limits_{h \to 0} \frac{{f\left( h \right) - 1}}{h} \hfill \\ f'\left( 0 \right) = \mathop {\lim }\limits_{h \to 0} \frac{{f\left( h \right) - f\left( 0 \right)}}{h} = \mathop {\lim }\limits_{h \to 0} \frac{{f\left( {h + 0} \right) - f\left( 0 \right)}}{h} = f\left( 0 \right)\mathop {\lim }\limits_{h \to 0} \frac{{f\left( h \right) - 1}}{h} \hfill \\ f'\left( 5 \right) = f\left( 5 \right)\frac{{f'\left( 0 \right)}}{{f\left( 0 \right)}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

with: \[f\left( 0 \right)=1\]

OpenStudy (mathmath333):

On differentiating the equation we get \(\large \color{black}{\begin{align} f(x+y)=f(x)f(y) \end{align}}\) with respect to x. \(\large \color{black}{ \begin{align} f'(x+y)\dfrac{d}{dx}(x+y)&=f'(x)f(y)+f'(y)f(x)\dfrac{dy}{dx}\\~\\ f'(x+y)(1+y')&= f'(x)f(y)+f(x)f'(y)(y') \\~\\ \end{align}}\) put \(\large x=0\) and \(\large y=5\) \(\large \color{black}{ \begin{align} f'(x+y)(1+y')&= f'(x)f(y)+f(x)f'(y)(y') \\~\\ f'(0+5)(1+5')&= f'(0)f(5)+f(0)f'(5)(5') \\~\\ f'(5)(1)&= f'(0)f(5)+f(0)f'(5)0 \\~\\ f'(5)(1)&= f'(0)f(5)\\~\\ f'(5)(1)&= 3\cdot 2\\~\\ \end{align}}\)

OpenStudy (anonymous):

Can we generalize and arrive at \(\dfrac{f'(x)}{f(x)}=3\)? We can tell that an exponential function will satisfy the required conditions. From here we can solve for \(f(x)\) and find \(f'(5)\).

OpenStudy (anonymous):

thank u